Prev Next

SW&IT Meeting Minutes 2021-02-22

Attendees: David, Igor, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210208
LR20210208 - Bec, Matthieu 72.0
LR20210208 - Edmeades, David 80.0
LR20210208 - Sola, Igor 17.5
LR20210208 - Grenz, Paul 40.0
LR20210208 - Hooper, Stephen 80.0
LR20210208 - Zhang, Yang 80.0
LR20210208 - Zhang, Xianyu 72.0

News from management

  • support calendar 2021A

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • SHARK NIR RFAs
    • review candidates, sw student
    • dms refactoring
  • This week
    • ATT tasks and priorities
    • dms refactoring, websockets
    • $ADOPT_LOG examination
    • Thu/Fri off

YZ

  • Previous:
    • Worked with Matthieu to set up the OSCO system.
    • Some TMS discussions (with eTalon engineers and within the TMS group).
    • Met with Paul for his TCS compilation test.
  • Next:
    • Debug AO-Cravat/OSCO system, and git/jenkins work.
    • Release TMS software and schedule a daytime test.
    • Follow-up with Olga for her LBC/robs request.

DE

  • Appointment moved to next week
  • Synology
    • RMA unit
      • Good shape, installed, working with support on details before setting time to swap unit
    • eBay unit
      • Major shipping damage
  • Cobbler: spun up VM, making notes for KS config
  • Backups
    • Google is ending unlimited storage for GSuite, moving to 100TB per institution total. This should be fine for us except for my offsite backups
    • Talking to Backblaze this week about said backups
  • PG interns working in lab, using wired Ethernet

XZ

  • Last week
    • ATT AdSec tasks and priorities
    • Optics/ASM for SHARK NIR
    • ASM set flat report script
    • AO support
  • This week
    • ATT AdSec and SW
    • AO SW ICD
    • AO support

PSG

  • Last week
    • Compiled changes to the OSS on tcs-test
  • This week
    • Update OSS GUI to reflect new swing arm control & white light source.

IS

Last week:
  • PIT release
    • changes to PEPSI, displaying correct status on manually modified proposal xml files, working on a encrypted key to display in addition to the proposal id.

This week:
  • PIT release
    • changes to PEPSI, displaying correct status on manually modified proposal xml files, working on a encrypted key to display in addition to the proposal id.
  • OT XSLT test cases
  • LUCI
    • Talk to Dave Thompson about dealing with hanging clients.
    • IT #7493 Aladin slow subtraction

SH

  • Technical support
  • E-mail list questions and maintenance
  • Dave T. robs complaints
  • Vacation May 8th
  • core-b down then up then down
  • Ilya and CJ and problems with connectivity from MPIA
  • Temporary key Steward
  • DIMM, though not too much.
  • RemoteEye II

next:

  • yumsrv -- rekick and reload to Fedora 32.
  • RPM maintenance
  • Network configuration git update
  • UITS
  • NAS/rsync AO maybe
Topic revision: r11 - 22 Feb 2021, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback